OpenStudy (anonymous):

So here is the question the sum of two numbers is 12 the difference between 3times of the first number and 5times of the second is 12 what are the numbers? could you teach it to me. step by step? pretty please? :>

OpenStudy (anonymous):

first no is 9 and second no is 3

OpenStudy (anonymous):

let x, y be the two no so according to ur question x+y=12 3x-5y=12 now solve the both equation and get the ans

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

x+y=12 x=12-y Using the second equation: 3(12-y)- 5y = 12 Solve the preceding equation for y. Plug that value of y into the first equation to solve for x.

OpenStudy (anonymous):

ohgord. i got it, thank you! so much

OpenStudy (anonymous):

Your welcome.